OpenStudy (anonymous):

Rewrite the following in Ax + By = C form: a. y=2x - 1 b. y=-9x - 5 c. y=6x - 6

OpenStudy (anonymous):

a) y-2x=-1 b) y+9x=-5 c) y-6x=-6

OpenStudy (anonymous):

sorry, use the communative property of addition to match your ax+by=c format.

jhonyy9 (jhonyy9):

2x-y=1 9x+y=-5 6x-y=6